OpenStudy (anonymous):

A store mixes red fescue worth $12 per pound and ryegrass worth $16 per pound. The mixture is to sell for $13 per pound, Find how much should be used to make a 504-pound mixture. How many pounds of red fescue should be used? How many pounds of ryegrass should be used?

OpenStudy (anonymous):

Hmm.. well, in a mixture, there is half and half of substances. So, I am thinking, you must do this what is 504 divided by 2?

OpenStudy (anonymous):

252

OpenStudy (anonymous):

Good. Then, you are correct! 252 pounds of ryegrass and 252 pounds of red fescue are going to be used.

OpenStudy (anonymous):

it says its wrong

OpenStudy (anonymous):

Hmm. That was what I thought was right. Let me et some help.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@AccessDenied May you help beth7878?

hero (hero):

f + r = 504 12f + 16r = 13*504